# Bring My Cup! Personalizing Vision-Language-Action Models with Visual Attentive Prompting

### v1

@@ Appendix E Discussion and Future Directions
VAP is a training-free framework that personalizes pre-trained Vision–Language–Action models for user-specific robotic manipulation by augmenting them with a modular front-end for personal-object recognition and instruction rewriting. Across two simulation benchmarks and a real-world robotic arm, VAP substantially improves instance-level success over non-personalized and token-based baselines. Our current instantiation is scoped to tabletop scenes with a single personalized object and short-horizon tasks, and it relies on off-the-shelf open-vocabulary detection and segmentation to provide accurate masks. When grounding is incorrect or inconsistent across camera views, these errors propagate to the visual prompts and can lead to failures (Appendix D). We view these aspects not as fundamental obstacles, but as concrete axes along which VAP can be strengthened—for example by incorporating stronger cross-view 3D grounding, lightly adapting the action head to better exploit visual prompts, and combining our module with planners that reason over sequences of personalized sub-tasks.
@@
Even at the level of object manipulation, there is substantial room to expand what personalization means. It is not only about which object to handle, but also how to handle it: one user may prefer that their mug is always grasped by the handle, that their laptop is never lifted by the screen, or that a fragile plant is moved only by supporting the pot from below. These preferences are naturally expressed as user-specific affordance maps over personal objects, conditioning not just the target but the contact locations, approach directions, and allowable forces. Beyond inanimate objects, users may ask robots to “take care of my cat” or “be gentle with my child’s favorite plush toy,” which requires combining personalization with safety- and comfort-aware interaction models. Extending VAP to learn such affordance-level preferences from a small number of demonstrations or corrections is a promising direction.
Looking further ahead, we see personalized manipulation as one component of a broader ecosystem of user-centered robots. Ideas from VAP could be used to personalize spatial conventions (e.g., learning what “my side of the desk” or “where my keys usually go” means for each user), to restore a preferred workspace layout after cleaning, or to execute multi-step routines such as “set up my desk for work” or “reset the living room for movie night”. As robots move into shared homes and offices, personalization may also need to reconcile competing preferences across multiple users and devices, and to generalize personal concepts across embodiments (phones, desktop agents, mobile robots) while retaining on-device, training-free adaptation for privacy and responsiveness. Incorporating additional modalities such as speech, prosody, gaze, touch, and even physiological signals could help capture subtler forms of intent, comfort, and trust. At the same time, any such system must respect boundaries around what counts as “personal”, provide transparency and control over stored references, and degrade gracefully when users update or revoke their preferences. We hope that VAP can serve as a simple mechanism for injecting user-specific signals into large VLAs, and that future work will build on this to personalize not only objects, but also concepts, routines, and long-term interactions in a safe and trustworthy way.

### v2

@@ Appendix G Discussion and Future Directions
VAP is a training-free framework that personalizes pre-trained Vision–Language–Action models for user-specific robotic manipulation by augmenting them with a modular front-end for personal-object recognition and instruction rewriting. Across two simulation benchmarks and a real-world robotic arm, VAP substantially improves instance-level success over non-personalized and token-based baselines. Our current instantiation is scoped to tabletop scenes with a single personalized object and short-horizon tasks, and it relies on off-the-shelf open-vocabulary detection and segmentation to provide accurate masks. When grounding is incorrect or inconsistent across camera views, these errors propagate to the visual prompts and can lead to failures (Appendix D). We view these aspects not as fundamental obstacles, but as concrete axes along which VAP can be strengthened—for example, by incorporating stronger cross-view 3D grounding, lightly adapting the action head to better exploit visual prompts, and combining our module with planners that reason over sequences of personalized sub-tasks.
@@
Even at the level of object manipulation, there is substantial room to expand the definition of personalization. It is not only about which object to handle, but also how to handle it: one user may prefer that their mug is always grasped by the handle, while another may require that a fragile plant is moved by supporting the pot from below. These preferences can be naturally modeled as user-specific affordance maps, conditioning not just the target selection but also contact locations and approach directions. Extending VAP to learn such fine-grained preferences from a small number of demonstrations is a promising direction for safer and more user-aligned manipulation.
Looking ahead, we envision personalization as a key enabler for long-term human-robot interaction. The principles of VAP could be adapted to personalize spatial concepts (e.g., learning what “my side of the desk” implies) or to execute context-aware routines such as “set up my workspace”. As robots operate in shared environments, future systems must also address the challenge of managing conflicting preferences across multiple users while ensuring data privacy. We hope that VAP serves as a foundational step toward injecting user-specific signals into large VLAs, paving the way for robotic agents that can adapt to individual needs, routines, and concepts in a safe and transparent manner.